Use the given information to find the number of degrees of​ freedom, the critical values chi Subscript Upper L Superscript 2χ2L and chi Subscript Upper R Superscript 2χ2R​, and the confidence interval estimate of sigmaσ. It is reasonable to assume that a simple random sample has been selected from a population with a normal distribution. Nicotine in menthol cigarettes 80​% ​confidence; n = 25​, s = 0.27 mg.

df = (Type a whole​ number.)

χ2L = ​(Round to three decimal places as​ needed.)

χ2R = ​(Round to three decimal places as​ needed.)

The confidence interval estimate of σ is __ mg < σ < __ mg. ​(Round to two decimal places as​ needed.)
Answer #1

Solution :

Given that,

s = 0.27

s2 = 0.0729

Degrees of freedom = df = n - 1 = 24

\chi2L = \chi 2\alpha/2,df = 33.196

\chi2R = \chi 21 - \alpha /2,df = 15.659

The 80% confidence interval for \sigma is,

\sqrt(n - 1)s2 / \chi 2\alpha/2 < \sigma < \sqrt (n - 1)s2 / \chi 21 - \alpha /2

  \sqrt24 * 0.0729 / 33.196 < \sigma < \sqrt 24 * 0.0729 / 15.659

0.23 mg < \sigma < 0.33 mg

(0.23 , 0.33)